# The Complexity of Arc-Connectedness Relation in the Plane

### v1

@@ 1. Introduction
for all $x,y\in X$, where $C(\mathbb{I},X)$ is the Polish space of continuous functions [^1] We note that the function $f$ in this definition is generally called a path rather than an arc, where the latter term is used to mean a homeomorphic image of $\mathbb{I}$. However, path-connectedness coincides with arc-connectedness in Hausdorff spaces (see [Ill25, Theorem 2.11]). from the unit interval $\mathbb{I}$ to $X$.
That any analytic equivalence relation on the Cantor space $2^{\mathbb{N}}$ can be realized as the arc-connectedness equivalence relation on a compact subset of $\mathbb{R}^{3}$ immediately follows from Theorem 4.1 of [Bec98], in which Becker asked how complex the arc-connectedness equivalence relation can be for compact subsets of $\mathbb{R}^{2}$. Then Becker and Pol showed that for any Polish subspace $X\subseteq\mathbb{R}^{2}$, the relation $\approx_{X}$ is Borel if and only if all arc components of $X$ are Borel [BP01]. In fact, it has been proven that this equivalence relation is Borel for any Polish subspace of $\mathbb{R}^{2}$ [DSR24].
@@
This upper bound turns out to be the best possible. One can actually embed the Cantor space into the plane suitably to see that $E_{0}$ is Borel reducible to the arc-connectedness relation on the Knaster continuum (also known as the bucket handle). Indeed, the same result is immediate from [Sol02] as composants of this continuum coincide with its arc components. The reader may consult [Kur68, §48, V-VI] for the construction and related continuum-theoretic properties of the Knaster continuum.
Acknowledgements. This paper is a complementary part of the author’s Master’s thesis [Uya25] which was written under the supervision of Burak Kaya and received partial financial support from TÜBİTAK BİDEB through the Program 2210-E. The author would like to thank Benjamin Vejnar for bringing the Knaster continuum and Solecki’s related work to his attention.
@@ 2. Borel Equivalence Relations
In this section, we recall the necessary definitions and results from the theory of Borel equivalence relations. Our reference for this section is [DJK94].
@@
We know that $E$ is hyperfinite if and only if it is the orbit equivalence of a Borel $\mathbb{Z}$-action on $X$. One can easily modify the proof of this fact to obtain the following slightly more general result (for example, see [KM04, Theorem 6.6] and the remark below it).
@@ 3. Upper Bound
In this section, we adopt some definitions, notations, and well-known facts from [DSR24] to prove the upper bound. Let $H\subseteq\mathbb{R}^{2}$ be a Polish subspace. We endow the set $\mathcal{K}(H)$ of all compact subsets of $H$ with the Vietoris topology, which is a Polish space. The set $\mathcal{J}(H)\subseteq\mathcal{K}(H)$ of all arcs in $H$ and the map $e:\mathcal{J}(H)\rightarrow\mathcal{K}(H)$ that assigns to each arc the set of its endpoints are Borel (see [BP01]). We define $e_{1}(J)=\min\left(e(J)\right)$ and $e_{2}(J)=\max\left(e(J)\right)$ for all $J\in\mathcal{J}(H)$ with respect to the lexicographic order on $\mathbb{R}^{2}$, which is Borel. The equivalence relation $\approx_{\mathcal{J}(H)}$ defined on $\mathcal{J}(H)$ of being in the same arc component is given by
@@
of the set $H$ was proven to be a Borel partition [DSR24].
@@
It is easy to show that $\varphi^{-1}(J)\subseteq K$ is a closed interval for any arc $J\in\mathcal{J}_{p,q}|_{C}$222with the only exception that if $C$ is a simple closed curve, then the inverse image of only one element of $\mathcal{J}_{p,q}|_{C}$ may violate this, in which case the subsequent observations and propositions can be easily modified. Any distinct arcs $J,J^{\prime}\in\mathcal{J}_{p,q}|_{C}$ can intersect only at their endpoints, and therefore $\varphi^{-1}(J)$ and $\varphi^{-1}(J^{\prime})$ are two intervals that possibly intersect at one of their endpoints (see [DSR24, Lemma 7]). This allows us to linearly order the elements of $\mathcal{J}_{p,q}|_{C}$. The following proposition shows that this ordering is, in fact, ”chain-like”.
@@
whose projection to $\mathcal{J}_{p,q}^{2}$ is the set, call $\mathcal{D}_{1}$, of pairs of disjoint consecutive arcs. Each $(J,J^{\prime})$-section of the Borel set is finite because $\Gamma$ is determined by the endpoints of $J$ and $J^{\prime}$. It follows from the Lusin-Novikov Theorem [Kec95, Theorem 18.10] that $\mathcal{D}_{1}$ is Borel. Second, we consider the Borel set
@@
Due to Moore’s Theorem [Pit70], $H_{1}$ has countably many $\approx_{H}$-classes, and so it is Borel reducible to the identity relation on $\mathbb{N}$. Lastly, $\approx_{H_{2}}$ is already the identity relation on $H_{2}$. Clearly, $\approx_{H_{1}}$ and $\approx_{H_{2}}$ are essentially hyperfinite. Then

### v2

@@ 1. Introduction
for all $x,y\in X$, where $C(\mathbb{I},X)$ is the Polish space of continuous functions [^1] We note that the function $f$ in this definition is generally called a path rather than an arc, where the latter term is used to mean a homeomorphic image of $\mathbb{I}$. However, path-connectedness coincides with arc-connectedness in Hausdorff spaces (see [illanes2025, Theorem 2.11]). from the unit interval $\mathbb{I}$ to $X$.
That any analytic equivalence relation on the Cantor space $2^{\mathbb{N}}$ can be realized as the arc-connectedness equivalence relation on a compact subset of $\mathbb{R}^{3}$ immediately follows from Theorem 4.1 of [becker1998], in which Becker asked how complex the arc-connectedness equivalence relation can be for compact subsets of $\mathbb{R}^{2}$. Then Becker and Pol showed that for any Polish subspace $X\subseteq\mathbb{R}^{2}$, the relation $\approx_{X}$ is Borel if and only if all arc components of $X$ are Borel [becker2001]. In fact, it has been proven that this equivalence relation is Borel for any Polish subspace of $\mathbb{R}^{2}$ [saintraymond2024].
@@
This upper bound turns out to be the best possible. One can actually embed the Cantor space into the plane suitably to see that $E_{0}$ is Borel reducible to the arc-connectedness relation on the Knaster continuum (also known as the bucket handle). Indeed, the same result is immediate from [solecki2002] as composants of this continuum coincide with its arc components. The reader may consult [kuratowski1968, §48, V-VI] for the construction and related continuum-theoretic properties of the Knaster continuum.
Related Work. We refer the reader to https://arxiv.org/abs/2601.00601, an extended version of this document written in collaboration with additional authors, for a more comprehensive analysis of connectedness relations on Polish spaces.
@@ 2. Borel Equivalence Relations
In this section, we recall the necessary definitions and results from the theory of Borel equivalence relations. Our reference for this section is [dougherty1994].
@@
We know that $E$ is hyperfinite if and only if it is the orbit equivalence of a Borel $\mathbb{Z}$-action on $X$. One can easily modify the proof of this fact to obtain the following slightly more general result (for example, see [kechris2004, Theorem 6.6] and the remark below it).
@@ 3. Upper Bound
In this section, we adopt some definitions, notations, and well-known facts from [saintraymond2024] to prove the upper bound. Let $H\subseteq\mathbb{R}^{2}$ be a Polish subspace. We endow the set $\mathcal{K}(H)$ of all compact subsets of $H$ with the Vietoris topology, which is a Polish space. The set $\mathcal{J}(H)\subseteq\mathcal{K}(H)$ of all arcs in $H$ and the map $e:\mathcal{J}(H)\rightarrow\mathcal{K}(H)$ that assigns to each arc the set of its endpoints are Borel (see [becker2001]). We define $e_{1}(J)=\min\left(e(J)\right)$ and $e_{2}(J)=\max\left(e(J)\right)$ for all $J\in\mathcal{J}(H)$ with respect to the lexicographic order on $\mathbb{R}^{2}$, which is Borel. The equivalence relation $\approx_{\mathcal{J}(H)}$ defined on $\mathcal{J}(H)$ of being in the same arc component is given by
@@
of the set $H$ was proven to be a Borel partition [saintraymond2024].
@@
It is easy to show that $\varphi^{-1}(J)\subseteq K$ is a closed interval for any arc $J\in\mathcal{J}_{p,q}|_{C}$222with the only exception that if $C$ is a simple closed curve, then the inverse image of only one element of $\mathcal{J}_{p,q}|_{C}$ may violate this, in which case the subsequent observations and propositions can be easily modified. Any distinct arcs $J,J^{\prime}\in\mathcal{J}_{p,q}|_{C}$ can intersect only at their endpoints, and therefore $\varphi^{-1}(J)$ and $\varphi^{-1}(J^{\prime})$ are two intervals that possibly intersect at one of their endpoints (see [saintraymond2024, Lemma 7]). This allows us to linearly order the elements of $\mathcal{J}_{p,q}|_{C}$. The following proposition shows that this ordering is, in fact, ”chain-like”.
@@
whose projection to $\mathcal{J}_{p,q}^{2}$ is the set, call $\mathcal{D}_{1}$, of pairs of disjoint consecutive arcs. Each $(J,J^{\prime})$-section of the Borel set is finite because $\Gamma$ is determined by the endpoints of $J$ and $J^{\prime}$. It follows from the Lusin-Novikov Theorem [kechris1995, Theorem 18.10] that $\mathcal{D}_{1}$ is Borel. Second, we consider the Borel set
@@
Due to Moore’s Theorem [pittman1970], $H_{1}$ has countably many $\approx_{H}$-classes, and so it is Borel reducible to the identity relation on $\mathbb{N}$. Lastly, $\approx_{H_{2}}$ is already the identity relation on $H_{2}$. Clearly, $\approx_{H_{1}}$ and $\approx_{H_{2}}$ are essentially hyperfinite. Then